OpenStudy (anonymous):

what is the least common multiple of 5 and 2? don't know bc one is even and the other is an odd number and I don't know why I cant reply????

OpenStudy (dangerousjesse):

Well, what's the prime factorization of 5?

OpenStudy (anonymous):

I have no idea to be honest?

OpenStudy (dangerousjesse):

Hint: if it can't be divided by a whole number, it's already prime :)

OpenStudy (dangerousjesse):

So 5=5 and 2=2 ^-^

OpenStudy (anonymous):

so 5 is a prime number?

OpenStudy (dangerousjesse):

Now can you find the largest power of 2 and 5 that appears in the prime factorizations?

OpenStudy (dangerousjesse):

Yes :)

OpenStudy (anonymous):

would that be 10?

OpenStudy (dangerousjesse):

Power= \(\huge x^y\)

OpenStudy (dangerousjesse):

So what's the largest y you can possibly have if x=5, to still have a prime number?

OpenStudy (dangerousjesse):

Hint: it's always the same answer :)

OpenStudy (anonymous):

I have no idea you lost me at xy

OpenStudy (dangerousjesse):

Do you know what 3 to the power of 2 equals?

OpenStudy (anonymous):

18 right

OpenStudy (dangerousjesse):

\(3\times 3=18?\)

OpenStudy (dangerousjesse):

What if I explained it like this: The "power" is just telling you how many times to multiply the number by itself. So, if you had a power of 4 and an integer of 10, you'd get \(10\times 10\times 10\times 10=10,000\)

OpenStudy (dangerousjesse):

It's just a way to shorten equations so you won't have any that say \("10\times 10\times 10\times 10\times 10\times 10\times 10\times 10\times 10"\)

OpenStudy (dangerousjesse):

So, when you multiply a number by itself more than once, it's no longer a prime number. Prime numbers \(cannot\) be divided by a whole number. So you have \(2^1\) and \(5^1\) as the highest power possible for the number to remain prime :)

OpenStudy (dangerousjesse):

When you multiply \(2^1\times 5^1\), you get an LCM of 10 :)